OpenStudy (anonymous):

Divide \[\frac{ -14y ^{5} }{ -7y ^{3} }\] Simplify your answer

OpenStudy (xapproachesinfinity):

Do you know this property \[\frac{ a^m }{ a^n }=a ^{m-n}\]

OpenStudy (xapproachesinfinity):

use that for this

OpenStudy (xapproachesinfinity):

\[\frac{ -14 }{ -7 }y ^{5-3}\]

OpenStudy (xapproachesinfinity):

continue...